7. Problems and Applications Q7 Consider the following data on U.S. GDP Nominal GDP GDP Deflator Year (Billions of dollars) (Base year 2009) 17,947 2015 109.8 1995 7,664 75.3 The growth rate of nominal GDP between 1995 and 2015 was and the growth rate of the GDP deflator between 1995 and 2015 was v. (Hint: The growth rate of a variable X over an year period is calculated as 100 x final initial Measured in 2009 prices, real GDP was billion in 1995 and billion in 2015. The growth rate of real GDP between 1995 and 2015 was The growth rate of nominal GDP between 1995 and 2015 was than the growth rate of real GDP
Answer #1

Real GDP = Nominal GDP/deflator x 100

The growth rate of nominal GDP between 1995 and 2015 was: 100 x ((17947/7664)1/20 -1) = 4.346%

The growth rate of GDP deflator between 1995 and 2015 was: 100 x ((109.8/75.3)1/20 -1) = 1.904%

Measured in 2009 prices:

Real GDP was : (7664/75.3) x 100 = $ 10177.95 bn in 1995

Real GDP was : (17947/109.8) x 100 = $ 16345.17 bn in 2015

The growth rate of real GDP between 1995 and 2015 was:  100 x ((16345.17/10177.95)1/20 -1) = 2.397%

The growth rate of nominal GDP between 1995 and 2015 was more than the growth rate of real GDP